OpenStudy (anonymous):

Please post answer so i can check my answer. Write the equation of the line that is parallel to the line 4x - 3y = -12 and passes through the point (-3, 4).

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

write it in the original form y = mx + b

OpenStudy (anonymous):

y=4/3x+4 but that is not the answer

Mehek (mehek14):

I know it isn't

Mehek (mehek14):

parallel lines have the same slope

Mehek (mehek14):

so the slope will remain 4/3

OpenStudy (anonymous):

okay. I finished the problem but I got y=4/3x+-4/3 and the answer choices are y=4/3x+8 and y=4/3x+3

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

ok so you have to plug in the points

OpenStudy (anonymous):

so 4=4/3(-3)+-4/3 ? @Mehek14

Mehek (mehek14):

yes solve it to find the y-intercept

OpenStudy (anonymous):

is the answer y=4/3x+8?

OpenStudy (anonymous):

@Mehek14

Mehek (mehek14):

yup Btw is this from flvs Algebra 1?

OpenStudy (anonymous):

yeah. why? @Mehek14

Mehek (mehek14):

from chapter 3?

OpenStudy (anonymous):

yeah @Mehek14

Mehek (mehek14):

I just finished ch3 and I'm like half way done for ch4